OpenStudy (anonymous):

To which subsets of the real numbers does the number √42 belong?

OpenStudy (anonymous):

It belongs to the irrational numbers.

OpenStudy (anonymous):

At first, you need to simplify this. To do that you can express 42 as a multiplication of primes.

OpenStudy (anonymous):

@Cheese3.1416 dont give answers right away.

OpenStudy (anonymous):

To which subsets of the real numbers does the number √42 belong? A. whole numbers, natural numbers, integers B. irrational numbers C. rational numbers D. whole numbers, integers, rational numbers

OpenStudy (anonymous):

Doing what I said, you get \[\sqrt{2\times 3\times 7}=\sqrt{42}\]. Can you see the answer from this?

OpenStudy (anonymous):

i have no clue im very very bad at math

OpenStudy (anonymous):

Do you know all of those subsets, or you are not sure of what all of them are?

OpenStudy (anonymous):

not sure what all of them are

OpenStudy (anonymous):

Ok, whole: 1, 2, 3, 4, 5, ... integers: ..., -3, -2, -1, 0, 1, 2, 3, ... rational: is a number that can be writen as a division of irrational: opposite from rational

OpenStudy (anonymous):

sorry, rational: is a number that can be writen as a division of integers

OpenStudy (anonymous):

so would the answer be B?

OpenStudy (anonymous):

im sorry if its wrong its been years since ive done math

OpenStudy (anonymous):

Wait, this is not trivial, no problem. It is known that sqrt of any prime is irrational, so the multiplication of sqrts of primes is also irrational. To do this, you'd have to know this, no problem if you didn't. So in this case the answer would be A

OpenStudy (anonymous):

sorry, B

OpenStudy (anonymous):

How did you first got to this answer?

OpenStudy (anonymous):

oh so i was right it is B? thanks for the help i understand alittle bit now i think im trying to get my high school diplma so i can go to college

OpenStudy (anonymous):

i got the answer when you explain that rational: is a number that can be writen as a division of integers

OpenStudy (anonymous):

Ok, but again, this is not trivial. Something like that might seem irrational but is actually rational: 5,7357357357357357357357357357357...=8372/9999. And that is even less intuitive for sqrts, so be carefull